OBJ.4 After the accounts are closed on February 3, prior to liquidating the partnershin. capital accounts of William Gerloff, Joshua Chu, and Courtney Jewett are $19,300, $4 and $22,300, respectively. Cash and noncash assets total $5,200 and $55,900, respecti Amounts owed to creditors total $15,000. The partners share inco of 2:1:1. Between February 3 and February 28, the noncash assets are sold for s the partner with the capital deficiency pays the deficiency to the partnershi PR 12-5A Statement of partnership liquidation P, the $4,500 0 ivel me and losses in the ratio 300 liabilities are paid.
ting for Partnerships and Limited Liability Companies 623 Instructions 1. Prepare a statement of partnership liquidation, indicating (a) the sale of a division of loss, (b) the payment of liabilities, (c) the receipt of the defi the appropriate partner), and (d) the distribution of cash sets and fromm 2. Assume that the partner with the capital to pay the deficiency. Journalize the entries to (a) allocate the partners deficiency and deficiency declares bankruptcy and is unable (b) distribute the remaining cash.
Answer #1

Answer:

1. Preparation of statement of partnership liquidation -

Particulars Cash Non Cash Capital Balances
Assets Creditors William Gerloff Joshua Chu Courtney Jewett
Preliquidation Balances $5,200 $55,900 $15,000 $19,300 $4,500 $22,300

(a). Sale of Non cash assets at a loss of $21,600 distributed at 2:1:1

34,300 (55,900) - (10,800) (5,400) (5,400)
Balances due 39,500 - 15,000 8,500 (900) 16,900
(b)Payment of liabilities (to creditors) (15,000) - (15,000) - - -
Balances due 24,500 - - 8,500 (900) 16,900
(c) Receipt of deficiency (from Joshua Chu) 900 - - - 900 -
Balances due 25,400 - - 8,500 - 16,900
(d)Distribution of Cash to the partners (25,400) - - (8,500) - (16,900)
Post liquidation balances - - - - - -

2. Journalizing the entries when the capital deficiency partner is declared as bankrupt-

(a). To allocate the partner's deficiency :

When a capital deficit partner is unable to pay his deficiency will shared by the other partners in their profit & loss sharing ratio ( here 2:1).

That is, Deficit of $ 900 will be shared as follows:

  • William Gerloff's share = 900 × 2/3 = $600
  • Courtney Jewett's share = 900 × 1/3 = $300

To record the allocation of partner's deficiency -

William Gerloff Capital $600
Courtney Jewett Capital $300

Joshua Chu Capital

$900

(b). To distribute the remaining cash :

Since the deficiency of Joshua was shared by William and Courtney , there was no receipt of cash. So, the cash balance after sale of assets and payment to creditors was $ 24,500. This is the cash remaining to be allocated to partner's in their existing capital balances ratio.

The capital balances of William Gerloff and Courtney Jewett after allocation of deficiency will be as follows:-

  • William Gerloff = 8,500 - 600 = $ 7,900
  • Courtney Jewett = 16,900 - 300 = $ 16,600

The cash balance of $ 24,500 will also be allocated in same ratio as above. i.e, $7,900 and $16,600 to William Gerloff and Courtney Jewett respectively.

To record the distribution of cash after liquidation -

William Gerloff Capital $7,900
Courtney Jewett Capital $16,600

Cash

$24,500